OpenStudy (anonymous):

Let f and g be functions that are neither even not odd. a. create an example where f+g is odd b. create an example where f times g is even c. create an example where f times g is odd

zepdrix (zepdrix):

Hmm here is something we could use for part a maybe.\[\Large\rm f=x^3+x^2,\qquad g=x^5-x^2\]Do you see what happens when we add f and g? What do we end up with?

OpenStudy (anonymous):

what about b and c @zepdrix